# Bob’s Confetti: Phonetic Memorization Attacks in Music and Video Generation

### v3

@@ 3.1 Motivation
3.2 Attack Prompt Construction
We formalize the two attack strategies—Adversarial PhoneTic Prompting (APT) and Adversarial VerbaTim Prompting (AVT)—by specifying the procedures used to construct their corresponding attack prompts.
Adversarial PhoneTic Prompting (APT).
Given a lyric sequence $L=\{w_{1},w_{2},\dots,w_{n}\}$, the objective of APT is to construct a modified sequence $L^{\prime}=\{w^{\prime}_{1},w^{\prime}_{2},\dots,w^{\prime}_{n}\}$ such that
$\Phi(w_{i},w^{\prime}_{i})\approx 1\qquad\forall i\in\{1,\dots,n\},$
where $\Phi(\cdot,\cdot)$ is a CMU Pronunciation Dictionary (CMUdict)-based (CMU, 1993) phonetic similarity function. CMUdict is a lexical resource that provides word-to-phoneme mappings, syllable counts, and stress patterns for over 100,000 English words, which make it particularly useful for tasks requiring phoneme-analysis such as rhyme detection or stress alignment.
We operationalize $\Phi$ as a vector of complementary phonetic similarity features:
$\Phi(a,b)=\big[\,S_{\mathrm{ph}}(a,b),\;S_{\mathrm{rh}}(a,b),\;S_{\mathrm{sy}}(a,b),\;S_{\mathrm{st}}(a,b),\;S_{\mathrm{jac}}(a,b),\;S_{\mathrm{cv}}(a,b),\;S_{\mathrm{vow}}(a,b)\,\big],$
here, $S_{\mathrm{ph}}(a,b)$ measures phoneme-sequence similarity using a SequenceMatcher ratio on CMUdict phoneme tokens, while $S_{\mathrm{rh}}(a,b)$ captures rhyme similarity through overlap of terminal phonemes. $S_{\mathrm{sy}}(a,b)$ encodes syllable-count matches derived from CMUdict vowel counts, and $S_{\mathrm{st}}(a,b)$ quantifies stress-pattern similarity based on alignment of CMUdict stress digits. To provide additional perspectives, $S_{\mathrm{jac}}(a,b)$ computes phoneme-level Jaccard similarity as a set overlap, $S_{\mathrm{cv}}(a,b)$ compares consonant–vowel (CV) patterns to reflect structural rhythm, and $S_{\mathrm{vow}}(a,b)$ evaluates vowel-core similarity by aligning stressed vowel phonemes.
Line-level similarity is computed by averaging word-level feature scores, and song-level similarity is obtained by averaging across lines. Figure 2 shows an example of modified lyrics of the famous Lose Yourself song generated using our APT attack.
Adversarial VerbaTim Prompting (AVT).
As a complementary upper bound, AVT directly reuses original verbatim lyrics. Formally, the input prompt is $L^{\prime}=L$, which forces the model to reproduce outputs highly overlapping with the original song. This attack isolates direct memorization pathways by measuring how closely generated outputs align with known copyrighted works.
4 Experiments
4.1 Setup
@@
We use SUNO and YuE to generate songs conditioned on both lyrics and genre descriptions. For APT attacks, we rely on SUNO, as it prevents users from generating songs with original verbatim lyrics. For AVT attacks, we evaluate both SUNO and YuE; however, in the case of SUNO, we were only able to produce songs in Mandarin or Cantonese, since—surprisingly—it did not flag any verbatim lyrics in those languages. We utilize Veo3 for video generation.
Automatic Evaluation (AudioJudge).
@@
Objective Metrics (MiRA).
We complement AudioJudge with two model-agnostic metrics from MiRA (Batlle-Roca et al., 2024): (i) CLAP similarity $s_{\text{CLAP}}\in[0,1]$, which measures high-level audio–text alignment, and (ii) CoverID $s_{\text{CID}}\in[0,1]$, which quantifies training-data overlap. Together, they capture memorization fidelity and replication likelihood. We independently verified to align strongly with human-rated judgments (Figure 7 of Appendix E), where we conduct manual listening tests. Participants rate similarity between original and generated clips on a 5-point Likert scale, explicitly instructed to ignore lexical content and focus on musical features (melody, rhythm, timbre). These human judgments provide a sanity check for automated metrics.
4.2 APT Prompt Generation
Genre Song (Artist) $S_{\mathrm{ph}}$ $S_{\mathrm{rh}}$ $S_{\mathrm{sy}}$ $S_{\mathrm{st}}$ $S_{\mathrm{jac}}$ $S_{\mathrm{cv}}$ $S_{\mathrm{vow}}$ $\Phi$
Rap HUMBLE (Kendrick Lamar) 0.622 0.624 0.925 0.943 0.612 0.739 0.645 0.730
DNA (Kendrick Lamar) 0.791 0.809 0.859 0.935 0.778 0.855 0.820 0.835
Lose Yourself (Eminem) 0.817 0.835 0.936 0.964 0.790 0.901 0.876 0.874
Pop Espresso (Sabrina Carpenter) 0.676 0.691 0.894 0.922 0.667 0.844 0.698 0.770
We Will Rock You (Queen) 0.630 0.636 0.840 0.868 0.610 0.863 0.640 0.727
Let It Be (The Beatles) 0.766 0.810 0.953 0.969 0.708 0.924 0.897 0.861
Korean APT (ROSÉ & Bruno Mars) 0.697 0.703 0.971 0.990 0.694 0.708 0.698 0.780
GENTLEMAN (PSY) 0.542 0.555 0.968 0.978 0.531 0.598 0.541 0.673
Christmas Jingle Bell (Traditional) 0.462 0.506 0.744 0.727 0.414 0.697 0.551 0.586
Jingle Bell Rock (Bobby Helms) 0.866 0.914 1.000 1.000 0.831 0.942 0.955 0.930
We assembled a candidate pool of $\approx 30$ songs from country-specific charts (U.S. Billboard Hot 100 and Korea Circle (Gaon), spanning decades, genres (rap, pop, ballad), and languages (English, Korean, Mandarin, Cantonese). Lyrics were normalized (case, punctuation, line breaks) to preserve cadence. For each song we synthesized three APT variants using Claude-3.5-Haiku under constraints that preserve phoneme sequence, rhyme, syllable count, and stress pattern (see Appendix G for the prompt). Candidates were scored based on the metric $\Phi$ and filtered at $\Phi\geq 0.65$ (further verified with human inspection). Because SUNO exposes no public API, we evaluated a stratified high-$\Phi$ subsample ($N\approx 30$); the final analysis uses $N=30$ APT and $N=16$ AVT generations chosen for high $\Phi$ and balanced coverage.
Table 1 shows that our APT rewrites reliably preserve meter: syllable-count ($S_{\mathrm{sy}}$) and stress-pattern ($S_{\mathrm{st}}$) scores are uniformly high across genres (many ≥ 0.90), indicating strong cadence preservation, while some show substantially lower $S_{\mathrm{ph}}$ and $S_{\mathrm{rh}}$. Phoneme Jaccard ($S_{\mathrm{jac}}$) and CV-pattern ($S_{\mathrm{cv}}$) largely track $S_{\mathrm{ph}}$, reinforcing that exact phoneme reuse and consonant–vowel structure co-occur where rewrites succeed. These patterns validate our $\Phi$-based filtering and explain why high-$\Phi$ candidates—particularly in rap and pop—are the strongest targets for phonetic-triggered memorization.
@@ 4.3 APT Attack Results
We evaluate the effectiveness and generality of our APT attack across a diverse set of songs, spanning genres (rap and pop), languages (English and Mandarin), and models (YuE and SUNO). We systematically modify lyrics to preserve phonetic structure—particularly rhyme and cadence—while discarding their original semantics. Our experiments demonstrate that such sub-lexical perturbations consistently elicit high-similarity outputs, revealing memorization behaviors that persist even under genre shifts, multilingual inputs, and model stochasticity. Results are grouped by musical domain to highlight trends and vulnerabilities specific to each category.
4.3.1 Iconic Songs (Rap)
Table 2: AudioJudge and MiRA similarity scores for phoneme-based and stylistic variations of “DNA” (Kendrick Lamar), “Lose Yourself” (Eminem), and “HUMBLE” (Kendrick Lamar). Melody, Rhythm, Harmony, Identity, and Style scores are derived from AudioJudge with gpt-4o-audio-preview. CLAP and CoverID are extracted from MiRA. All samples were generated using SUNO.
Song (Artist) Genre Variant AudioJudge MiRA
DNA (Kendrick Lamar) rap (gen1) 0.90 0.95 0.95 0.98 0.96 0.699 0.183
rap (gen2) 0.90 0.95 0.90 0.80 0.92 0.659 0.343
rap (gen3) 0.90 0.85 0.85 0.92 0.90 0.664 0.175
gangsta, rap, trap 0.70 0.85 0.90 0.85 0.92 0.687 0.219
Lose Yourself (Eminem) intense rap 0.80 0.85 0.95 0.60 0.80 0.773 0.147
N/A 0.70 0.65 0.95 0.70 0.80 0.683 0.255
HUMBLE (Kendrick Lamar) trap gangsta, bold, sparse, direct 0.95 0.97 0.95 0.98 0.96 0.740 0.160
rap gangsta, bold, sparse, direct 0.90 0.95 0.90 0.95 0.92 0.725 0.190
Table 2 reports results for three iconic rap tracks generated with APT attack: “DNA” (Kendrick Lamar), “Lose Yourself” (Eminem), and “HUMBLE” (Kendrick Lamar). Despite these alterations, the generated outputs consistently preserved core musical attributes such as melody, rhythm, and harmony, demonstrating the resilience of phonetic mimicry in guiding model behavior.
For “DNA”, all phoneme-based generations achieved strong melodic and rhythmic fidelity (0.90 and 0.85–0.95, respectively), with harmony and style also remaining robust. Even under gangsta/trap conditioning, performance stayed high, and MiRA confirmed this trend with moderate CLAP similarity (≈0.66–0.70) and low-to-mid CoverID values (0.17–0.34). “Lose Yourself” showed a similar pattern: phoneme substitutions preserved cadence, with the “intense rap” variant achieving melody 0.80 and rhythm 0.85 alongside high harmony (0.95); CLAP rose to 0.77 while CoverID stayed relatively low (0.15–0.25). “HUMBLE” produced the strongest results, with trap- and rap-styled variants nearing baseline quality (melody: 0.95, rhythm: 0.97, identity: 0.98), corroborated by CLAP ≈0.74 and low CoverID (0.16–0.19).
Taken together, these results reveal a consistent vulnerability in lyrics-to-song generation: phoneme-preserving distortions yield high-fidelity outputs that rival or exceed those produced with original lyrics. AudioJudge scores confirm strong alignment across melody, rhythm, and harmony, while MiRA metrics further show that these variants remain musically similar (high CLAP) without being flagged as direct replicas (low CoverID). This exposes a sub-lexical weakness where rhyme, rhythm, and phonetic shape dominate over semantics, enabling unintended memorization leakage.
4.3.2 Iconic Songs (Pop)
Table 3: AudioJudge and MiRA similarity scores for English and Korean iconic song recreations from lyrics using SUNO. Melody, Rhythm, Harmony, Identity, and Style scores are from AudioJudge (gpt-4o-audio-preview), while CLAP and CoverID are from MiRA. All songs were generated with no genre description provided.
Language Song (Artist) AudioJudge MiRA
English Espresso (Sabrina Carpenter) 0.90 0.95 0.80 0.95 0.88 0.829 0.105
Let It Be (The Beatles) 0.90 0.85 0.90 0.60 0.75 0.639 0.349
Can’t Help Falling in Love (Elvis Presley) 0.95 0.85 0.90 0.60 0.75 0.551 0.405
We Will Rock You (Queen) 0.85 0.90 0.80 0.40 0.75 0.518 0.423
Korean APT (ROSÉ & Bruno Mars) 0.95 0.98 0.80 0.75 0.88 0.852 0.119
Gangnam Style (PSY) 0.95 0.97 0.95 0.85 0.96 0.801 0.210
GENTLEMAN (PSY) 0.85 0.90 0.80 0.95 0.88 0.830 0.334
We tested iconic English and Korean pop tracks on two representative L2S systems: the commercial black-box model SUNO and the open-source model YuE (Table 3). Each track was paired with APT-attacked lyrics that preserved prosody while distorting semantics, yielding a multilingual benchmark for probing sub-lexical memorization vulnerabilities.
SUNO shows strong susceptibility across diverse songs. For APT (ROSÉ and Bruno Mars), adversarial lyrics reached near-perfect fidelity (melody: 0.95, rhythm: 0.98; CLAP: 0.852, CoverID: 0.119). Other tracks, including Espresso, Gangnam Style, and Let It Be, likewise preserved melodic and rhythmic alignment (0.85–0.97) despite distorted semantics, with moderate CLAP scores (0.64–0.83) and generally low CoverID (0.10–0.35). Classic ballads such as Can’t Help Falling in Love and We Will Rock You maintained strong cadence (melody: 0.85–0.95, rhythm: 0.85–0.90) though CoverID values rose above 0.40, indicating closer resemblance to training data. These results highlight that rhyme and rhythm, rather than meaning, dominate SUNO’s generations (See Table 6 of Appendix C.1 for additional APT attack results on Christmas Songs).
Taken together, these results demonstrate that phoneme-preserving substitutions consistently preserve musical elements across systems, languages, and genres. AudioJudge confirms robust alignment on melody, rhythm, and style, while MiRA reveals that these adversarial variants remain musically similar without always being direct replicas—exposing how L2S models rely heavily on phonetic rhythm and rhyme, posing clear memorization and copyright risks.
4.3.3 APT Attack against Veo 3
Given the success of our attack in L2S models, we next investigated how APT extends to lyrics-conditioned text-to-video (T2V) generation, where we conducted a case study using Veo 3 DeepMind (2024a; b), a recent multimodal video synthesis model. Here, the goal is to generate human speech in addition to other accompanying modalities (background music, video frames), conditioned on the transcript of the target generation. We evaluate whether phonetic cues alone — without explicit visual or semantic guidance — could trigger memorized visual outputs. Prompts were submitted using Veo 3’s "transcript mode" with only a minimal instruction prepended: "video with the following transcript:" followed by the respective lyrics.
In the Lose Yourself generations using APT attack, Veo 3 consistently produced a male rapper wearing a hoodie (Figure 4 of Appendix B). Notably, the output voice was rhythmically well-aligned with the original track, despite no mention of gender, clothing, setting, or musical style in the prompt. Similarly, for Jingle Bells, even with heavily phoneme-altered lyrics, the generated music retained the original song’s melody and rhythmic phrasing, underscoring the model’s reliance on phonetic rhythm as a cue for memorization.
These findings suggest that even phonetically similar but semantically meaningless prompts can trigger the reconstruction of memorized visual motifs. This extends beyond prior demonstrations of text-to-image or audio-only memorization and highlights a new axis of risk in generative multimodal models. While Veo 3 showcases remarkable video coherence, it also appears susceptible to sub-perceptual prompt leakage: a subtle but powerful form of memorization where phoneme patterns alone act as implicit keys to stored training examples. These results further underscore the need for dedicated memorization audits in text-to-video and lyrics-to-video systems, especially as such tools become increasingly integrated into creative pipelines. Future work should explore whether this behavior arises from overrepresentation of iconic music videos in the training distribution, and how phonetic conditioning interacts with visual token generation.

### v4

@@ 4.2.3 Adversarial VerbaTim Prompting (AVT)
5 Empirical Evaluation
5.1 Experimental Setup
@@
We use SUNO [sunoai] and YuE [yuan2025yue] to generate songs conditioned on both lyrics and genre descriptions. For APT attacks, we rely on SUNO, as it prevents users from generating songs with original verbatim lyrics. For AVT attacks, we evaluate on both SUNO and YuE; however, in the case of SUNO, we were only able to produce songs in Mandarin or Cantonese, since it did not flag any verbatim lyrics in those languages. We utilize Veo 3 for video generation.
Automatic Evaluation (AudioJudge)
We use AudioJudge [manakul2025audiojudge], with a gpt-4o-audio-preview backbone model. For each original–generated pair (x,x^)($\hat{x}$), AudioJudge assigns similarity scores across five dimensions: melody, rhythm, harmony, identity, and style:
$S_{\text{AJ}}(x,\hat{x})=\big(s_{\text{mel}},s_{\text{rhy}},s_{\text{har}},s_{\text{id}},s_{\text{sty}}\big)\in[0,1]^{5}.$
We validate AudioJudge’s discriminative capability in Section 6.1, confirming it reliably distinguishes matched from mismatched pairs. The system prompt and example outputs are provided in Figure 13 of Appendix I.
Objective Evaluation (MiRA)
We complement AudioJudge with two model-agnostic metrics from MiRA [batlle2024towards]: (i) CLAP similarity $s_{\text{CLAP}}\in[0,1]$, which measures high-level audio–text alignment, and (ii) CoverID $s_{\text{CID}}\in[0,1]$, which quantifies training-data overlap. Together, they capture memorization fidelity and replication likelihood. We independently verified that these metrics align strongly with human-rated judgments (Figure 7), where we conduct manual listening tests. Participants rate similarity between original and generated clips on a 5-point Likert scale, explicitly instructed to ignore lexical content and focus on musical features (melody, rhythm, timbre). These human judgments provide a sanity check for automated metrics.
5.2 APT Prompt Generation
We assembled a candidate pool of $\approx 30$ songs from country-specific charts (U.S. Billboard Hot 100 and Korea Circle (Gaon)), spanning decades and genres (rap, pop, ballad). For APT, we focus on two languages—English and Korean—whose typologically distinct phonological systems (stress-timed vs. syllable-timed) provide a rigorous test of generalizability. For AVT, we additionally include Mandarin and Cantonese to probe memorization under verbatim prompting across four languages total. Lyrics were normalized (case, punctuation, line breaks) to preserve cadence. For each song we synthesized three APT variants using Claude-3.5-Haiku under constraints that preserve phoneme sequence, rhyme, syllable count, and stress pattern (see Figure 12 of Appendix F for prompt). Candidates were scored based on the metric $\Phi$ and filtered at $\Phi\geq 0.65$ (further verified with human inspection). Because SUNO has no public API, we evaluated a stratified high-$\Phi$ subsample ($N\approx 30$); the final analysis uses $N=30$ APT and $N=16$ AVT generations chosen for high $\Phi$ and balanced coverage. Figure 3 shows an example of modified lyrics of the famous Lose Yourself song generated using our APT attack.
Table 1 shows that our APT rewrites reliably preserve meter: syllable-count ($S_{\mathrm{sy}}$) and stress-pattern ($S_{\mathrm{st}}$) scores are uniformly high across genres (many ≥ 0.90), indicating strong cadence preservation, while some show substantially lower $S_{\mathrm{ph}}$ and $S_{\mathrm{rh}}$. Phoneme Jaccard ($S_{\mathrm{jac}}$) and CV-pattern ($S_{\mathrm{cv}}$) largely track $S_{\mathrm{ph}}$, reinforcing that exact phoneme reuse and consonant–vowel structure co-occur where rewrites succeed. These patterns validate our $\Phi$-based filtering and explain why high-$\Phi$ candidates (particularly in rap and pop) are the strongest targets for phonetic-triggered memorization.
5.3 Baseline Validation
Table 2: Baseline Comparison. AudioJudge similarity scores (mean $\pm$ std) for different input conditions. APT achieves significantly higher similarity across all musical dimensions compared to baselines.
Condition Melody Rhythm Harmony
Random Lyrics $0.10\pm 0.04$ $0.19\pm 0.06$ $0.12\pm 0.05$
Semantic Paraphrase $0.60\pm 0.12$ $0.13\pm 0.09$ $0.54\pm 0.14$
@@ 5.3 Baseline Validation
To isolate the role of phonetic structure in triggering memorization, we evaluate two control conditions. (1) Random Lyrics: We generated nonsensical lyrics from phonetically dissimilar word pools (e.g., “umbrella,” “xylophone,” “refrigerator”), randomly combined with no connection to original lyrics while matching structural format (line count, verse groupings). We generated 6 samples via SUNO (4 for Lose Yourself, 2 for DNA). (2) Semantic Paraphrase: We manually crafted meaning-preserving rewrites that deliberately break phonetic structure—using different vocabulary, syllable counts, and avoiding rhyme preservation. For example, “His palms are sweaty, knees weak, arms are heavy” becomes “He’s nervous and trembling, his limbs feel like lead weights.” We generated 8 samples via SUNO across 4 songs.
Table 3: APT Attack Results. AudioJudge similarity scores (mean $\pm$ std) and MiRA metrics for rap, Billboard pop, and K-pop song recreations generated using SUNO. Rap songs: HUMBLE, DNA (Kendrick Lamar); Lose Yourself (Eminem). Billboard pop songs: We Will Rock You (Queen); Espresso (Sabrina Carpenter); Can’t Help Falling in Love (Elvis Presley); Let It Be (The Beatles). K-pop songs: GENTLEMAN, Gangnam Style (PSY); APT (ROSÉ & Bruno Mars); DOPE (BTS). Rap songs were generated using the genre variant “rap”, while Billboard pop and K-pop songs were generated directly from lyrics with no genre description provided. Melody, Rhythm, Harmony, Identity, and Style are from AudioJudge. CLAP and CoverID are from MiRA.
Category Song AudioJudge MiRA
Rap HUMBLE 0.90 $\pm$ 0.04 0.93 $\pm$ 0.03 0.88 $\pm$ 0.02 0.77 $\pm$ 0.13 0.92 $\pm$ 0.01 0.71 $\pm$ 0.13 0.12 $\pm$ 0.04
Lose Yourself 0.56 $\pm$ 0.21 0.73 $\pm$ 0.15 0.59 $\pm$ 0.24 0.48 $\pm$ 0.22 0.65 $\pm$ 0.24 0.63 $\pm$ 0.07 0.30 $\pm$ 0.15
DNA 0.85 $\pm$ 0.09 0.90 $\pm$ 0.05 0.90 $\pm$ 0.04 0.89 $\pm$ 0.07 0.93 $\pm$ 0.02 0.70 $\pm$ 0.02 0.23 $\pm$ 0.08
Billboard Pop We Will Rock You 0.60 $\pm$ 0.20 0.48 $\pm$ 0.36 0.52 $\pm$ 0.29 0.33 $\pm$ 0.23 0.51 $\pm$ 0.32 0.59 $\pm$ 0.05 0.30 $\pm$ 0.07
Espresso 0.43 $\pm$ 0.05 0.51 $\pm$ 0.14 0.45 $\pm$ 0.08 0.29 $\pm$ 0.03 0.51 $\pm$ 0.17 0.66 $\pm$ 0.03 0.12 $\pm$ 0.02
Can’t Help Falling in Love 0.86 $\pm$ 0.09 0.85 $\pm$ 0.07 0.89 $\pm$ 0.06 0.59 $\pm$ 0.09 0.84 $\pm$ 0.11 0.57 $\pm$ 0.08 0.49 $\pm$ 0.29
Let It Be 0.45 $\pm$ 0.13 0.53 $\pm$ 0.12 0.38 $\pm$ 0.16 0.35 $\pm$ 0.09 0.52 $\pm$ 0.15 0.64 $\pm$ 0.01 0.49 $\pm$ 0.12
K-pop GENTLEMAN 0.93 $\pm$ 0.03 0.97 $\pm$ 0.02 0.94 $\pm$ 0.02 0.89 $\pm$ 0.04 0.95 $\pm$ 0.02 0.73 $\pm$ 0.10 0.32 $\pm$ 0.11
Gangnam Style 0.90 $\pm$ 0.00 0.94 $\pm$ 0.01 0.90 $\pm$ 0.03 0.75 $\pm$ 0.00 0.92 $\pm$ 0.01 0.75 $\pm$ 0.12 0.13 $\pm$ 0.01
APT 0.89 $\pm$ 0.03 0.93 $\pm$ 0.02 0.90 $\pm$ 0.03 0.79 $\pm$ 0.06 0.92 $\pm$ 0.01 0.80 $\pm$ 0.05 0.15 $\pm$ 0.03
DOPE 0.80 $\pm$ 0.09 0.87 $\pm$ 0.05 0.85 $\pm$ 0.03 0.73 $\pm$ 0.10 0.88 $\pm$ 0.06 0.73 $\pm$ 0.01 0.20 $\pm$ 0.05
Table 2 presents the results. Random lyrics achieve negligible similarity (average: 0.14), while semantic paraphrases show moderate alignment (average: 0.42). APT prompts, by contrast, achieve consistently high scores (average: 0.91). This ordering confirms our central hypothesis: sub-lexical acoustic structure—not lexical content or semantic meaning—serves as the retrieval key for memorized training data. MiRA metrics provide complementary evidence: high CLAP scores (0.570.570.800.80) indicate strong audio-text alignment, while moderate CoverID values (0.120.120.490.49) suggest the outputs are musically similar without being exact replicas—precisely the regime where copyright concerns are most acute.
@@ 5.4 APT Attack Results
We evaluate the effectiveness of our APT attack across 11 songs spanning three categories: rap, Billboard pop, and K-pop. For each song, we generate 3 variants using SUNO and report mean $\pm$ std across generations. Table 3 and Table 8 of Appendix G presents the full results.
Rap Songs.
Rap tracks exhibit strong susceptibility to phonetic memorization due to their tight coupling between lyrics and rhythmic structure. HUMBLE (Kendrick Lamar) achieves the highest fidelity (melody: $0.90\pm 0.04$, rhythm: $0.93\pm 0.03$), with consistently low variance across generations. DNA shows similar robustness (melody: $0.85\pm 0.09$, rhythm: $0.90\pm 0.05$), confirming that phoneme-preserving prompts reliably trigger memorized outputs. Notably, Lose Yourself (Eminem) shows higher variance (melody: $0.56\pm 0.21$), suggesting that some songs require higher phonetic fidelity to trigger consistent memorization. Despite this variance, MiRA metrics remain strong (CLAP: $0.63\pm 0.07$), indicating underlying musical similarity even when AudioJudge scores fluctuate.
Billboard Pop.
Pop songs show more heterogeneous results. Classic tracks with distinctive melodic signatures—such as Can’t Help Falling in Love (Elvis Presley)—achieve high similarity (melody: $0.86\pm 0.09$, rhythm: $0.85\pm 0.07$). In contrast, songs with more complex harmonic structures like Let It Be show moderate scores (melody: $0.45\pm 0.13$), though still substantially above random baselines (Section 5.3). We Will Rock You and Espresso exhibit the lowest scores in this category, potentially due to their reliance on production elements (stomping rhythm, synth textures) that phonetic mimicry alone cannot capture.
K-pop.
K-pop tracks demonstrate the strongest and most consistent memorization across our evaluation. GENTLEMAN (PSY) achieves near-perfect scores (melody: $0.93\pm 0.03$, rhythm: $0.97\pm 0.02$), with minimal variance indicating highly deterministic memorization behavior. Gangnam Style and APT (ROSÉ & Bruno Mars) show similarly strong results, with all metrics exceeding 0.89. The consistently high scores and low variance in K-pop suggest that SUNO may have particularly strong memorization of these globally popular tracks, or that the phonetic structure of Korean lyrics provides especially effective retrieval cues.